Solve
(D^4+2D^3-3D^2)y=x^2+3e^2x+4sinx
1
Expert's answer
2020-10-12T18:37:02-0400

Auxiliary equation is "D^4 +2D^3-3D^2 = 0" (1)

Solving it, we get "D^2(D-1)(D+3) = 0"

"D = 0,0,1,-3"

Then "y = c_1+c_2x+c_3e^{x}+c_4e^{-3x}"


For particular integral,

For "x^2",

Let "y = ax^4 + bx^3+cx^2 +dx+e" be it's solution

then

"y'' = 12ax^2 +6bx +2c"

"y''' = 24ax + 6b"

"y'''' = 24a"

Putting values in equation "D^4+2D^3-3D^2 = x^2" , we get

"-36ax^2 + (48a-6b)x + (24a+12b-6c) = x^2"

comparing powers,

we get "a = -\frac{1}{36}, b = -\frac{2}{27}, c = -\frac{7}{27}"

then P.I. for "x^2" is, "-\frac{1}{36}x^4 - \frac{2}{27}x^3 - \frac{7}{27}x^2"




P.I. for "3e^{2x}" ,

"\frac{1}{D^4+2D^3-3D^2}e^{2x} = e^{2x}\frac{1}{2^4+(2\times 2^3)-(3\times 2^2)} = \frac{3}{20}e^{2x}"



P.I. for 4sin(x)

"\frac{1}{D^4+2D^3-3D^2}4sin(x) = \frac{1}{D^2(D^2+2D-3)}4sin(x)"

"= \frac{1}{(-1^2)(-1^2+2D-3)}4sin(x) = \frac{1}{2-D}2sin(x) = -2\frac{(D+2)}{D^2 - 4}sin(x)"

"= -2\frac{cos(x)+2sin(x)}{-1-4} = \frac{2}{5}(cos(x)+2sin(x))"


Solution of the differential equation is

"y = c_1+c_2x+c_3e^{x}+c_4e^{-3x}-\frac{1}{36}x^4 - \frac{2}{27}x^3 - \frac{7}{27}x^2+\frac{3}{20}e^{2x} +\frac{2}{5}(cos(x)+2sin(x))"